# Continuous Audio Language Models

### v2

@@ Abstract
Audio Language Models (ALM) have emerged as the dominant paradigm for speech and music generation by representing audio as sequences of discrete tokens. Yet, unlike text tokens, which are invertible, audio tokens are extracted from lossy codecs with a limited bitrate. As a consequence, increasing audio quality requires generating more tokens, which imposes a trade-off between fidelity and computational cost. We address this issue by studying Continuous Audio Language Models (CALM). These models instantiate a large Transformer backbone that produces a contextual embedding at every timestep. This sequential information then conditions an MLP that generates the next continuous frame of an audio VAE through consistency modeling. By avoiding lossy compression, CALM achieves higher quality at lower computational cost than their discrete counterpart. Experiments on speech and music demonstrate improved efficiency and fidelity over state-of-the-art discrete audio language models, facilitating lightweight, high-quality audio generation. Samples are available at hf.co/spaces/kyutai/calm-samples.
@@ 1 Introduction
This motivates an alternative strategy: autoregressive modeling of continuous latents without quantization. Standard variational autoencoders (VAEs) are easier to train, are not affected by issues such as codebook collapse, and can reconstruct audio at higher fidelity for the same latent dimensionality. Pioneering work in the vision domain that autoregressively models continuous sequences includes GIVT (Tschannen et al., 2024) and MAR (Li et al., 2024), followed by some larger models in the image domain (Fan et al., 2025; Gu et al., 2025) and attempts in the audio domain (Turetzky et al., 2024; Pasini et al., 2024b; Jia et al., 2025). In MAR, the authors model the per-token probability distribution with a diffusion model (in the form of a small MLP) conditioned on a latent variable modeled by an autoregressive transformer backbone.
SALAD (Turetzky et al., 2024) and DiTAR (Jia et al., 2025) adapt MAR-style diffusion heads for text-to-speech modeling, achieving better audio quality than discrete baselines. However, these works are limited to text-to-speech on small-scale and domain-specific datasets, leaving open the question of how well they can adapt to richer audio domains such as music.
@@
In this work, we propose Continuous Audio Language Models (CALM) that predict sequences in the latent space of a VAE, bypassing the need for quantization. While we build on the MAR architecture where a transformer backbone uses $(\mathbf{x}^{1},\ldots,\mathbf{x}^{s-1})$ to predict an intermediate latent $\mathbf{z}^{s}$, which then conditions a head (diffusion model) that models $p(\mathbf{x}^{s}|\mathbf{z}^{s})$, we find that without further improvements, it fails to generate rich audio content. To overcome this, we introduce several key innovations:
@@
1. Improving quality and stability: To mitigate error accumulation during inference, we follow Pasini et al. (2024b) and, during training, inject noise into the long-term context $(\mathbf{x}^{1},\ldots,\mathbf{x}^{s-1})$. Additionally, we introduce a short-context transformer that summarizes recent clean latents, providing the sampling head with both coarse long-range context and fine-grained local information. 2. Diffusion-to-Consistency replacement: We replace the diffusion model with a continuous consistency model (Lu & Song, 2025) during training, significantly accelerating inference without compromising sample quality. This change reduces the inference time of the sampler head by a factor of up to $\times 20$ in our music experiments and $\times 12$ in our speech experiments compared to the use of a model that uses an RQ-Transformer as head. This method is modality-agnostic and may benefit other domains such as image generation. 3. Gaussian Temperature sampling: Temperature control is crucial for high-quality speech generation, yet consistency models lack a formal mechanism for temperature sampling. We present a heuristic that approximates temperature sampling in the consistency setup. 4. Head batch multiplier: Sampling multiple noise levels at training time for the same latent highly accelerates training for a small cost.
Our model generates high-quality speech and music with lower inference time, outperforming state-of-the-art discrete-token autoregressive models in both efficiency and audio fidelity. To our knowledge, this is the first application of autoregressive continuous latent modeling to both domains.
@@ 2 Related Work
Autoregressive audio language models. Early autoregressive audio models operated on raw waveforms such as WaveNet (van den Oord et al., 2016) or learned discrete codes via VQ-VAE, such as in Jukebox (Dhariwal et al., 2020). The advent of neural audio codecs (Zeghidour et al., 2021; Défossez et al., 2024a) enabled high-quality vector-quantized tokenizers for general audio. These, in turn, powered discrete-token audio LMs: AudioLM (Borsos et al., 2023) for unconditioned audio generation while AudioGen (Kreuk et al., 2023), MusicLM (Agostinelli et al., 2023), and MusicGen (Copet et al., 2023) apply similar methods for text-to-audio and text-to-music generation. In speech, autoregressive modeling of RVQ codes has been used for text-to-speech generation (Wang et al., 2023; Kharitonov et al., 2023) as well as for spoken dialogue (Défossez et al., 2024b) or translation (Labiausse et al., 2025). However, all of these systems rely on lossy quantization, which inevitably degrades audio fidelity unless a significant compute budget is spent to generate a deep hierarchy of RVQ tokens. This is unlike CALM, which predicts continuous embeddings in one pass, providing a better quality-computation trade-off.
@@
Continuous autoregressive models. In GIVT, Tschannen et al. (2024) use a transformer to autoregressively model the latent space of a VAE trained on images, parameterize a Gaussian Mixture Model, and train it with a cross-entropy loss. The authors of MAR (Li et al., 2024) obtain better results by replacing the Gaussian mixture model with a diffusion-model enabling the approximation of more flexible distributions. In MAR, a large transformer backbone predicts a continuous embedding $\mathbf{z}^{s}$ given $(\mathbf{x}^{1},\ldots,\mathbf{x}^{s-1})$, which then conditions a small MLP diffusion network that models the probability distribution $p(\mathbf{x}^{s}|\mathbf{z}^{s})$ of the next latent. This eliminates the need for discrete tokenizers, but at the cost of slow sampling: MAR typically needs hundreds of denoising steps per token. Hang et al. (2025) aims to speed up MAR by replacing the diffusion head with a shortcut head (Frans et al., 2025) for few-step sampling. Shortcut models combine a diffusion loss and a self-consistency loss in order to accelerate the diffusion process. They reduce the number of diffusion steps from 100 to 8 with a similar image quality. Apart from this work, we are not aware of other works that try to reduce the inference time of the diffusion process in continuous autoregressive models. Remarkably, CALM achieves a quality comparable to the best discrete models with only one step of consistency modeling.
@@
In audio, the approach of MAR has been adapted for the task of Text-to-Speech (TTS) synthesis. SALAD (Turetzky et al., 2024) introduces a zero-shot TTS model that operates on continuous speech representations using a per-token latent diffusion process. By leveraging semantic tokens for contextual information and determining synthesis stopping points, SALAD achieves improved intelligibility and audio quality without relying on quantization. Similarly, DiTAR (Jia et al., 2025) presents a patch-based autoregressive framework combining a language model with a diffusion transformer for speech generation. This approach models dependencies between aggregated local patches of continuous tokens, using a causal language model to produce embeddings, which, along with previous patches, serve as inputs to a bidirectional diffusion transformer that predicts the next patch. The authors observe that providing local context through patching was determinant to improve their model, which corroborates what we observe with the introduction of a short context transformer into our model. While these two studies focus on TTS modeling, we extend the method to music. In Pasini et al. (2024b), the authors apply the MAR framework to music generation using a relatively small dataset comprising 20,000 single-instrument stems, training their model on 10-second excerpts on top of their continuous compression model Music2Latent (Pasini et al., 2024a). They introduce a method for noise augmentation of the data that allows the model to avoid error accumulation. However, we notice that scaling to a more complex and diverse dataset consisting of full musical pieces makes their approach struggle to maintain high-quality generation on longer sequences. To address these challenges, we propose novel strategies that enhance generation quality while also improving inference efficiency. Finally, Music2Latent2 (Pasini et al., 2025) explores combining autoregressive and consistency modeling but for compression.
@@ 3 Background
This approach enables efficient parallel modeling of RVQ sequences, allowing all codebooks corresponding to a single timestep to be generated simultaneously without introducing any delay. However, a key limitation lies in the use of the RQ-Transformer, which is computationally intensive; as its resource requirements scale with the number of RVQ.
@@ 3.2 Consistency Models
By using $T=\frac{\pi}{2}$ and $\alpha_{t}=\cos(t),\sigma_{t}=\sin(t)$, Lu & Song (2025) derive the following continuous-time consistency loss where $w_{\psi}(t)$ is an adaptive weighting function:
@@ 4.1 Our VAE-GAN
While training the model with adversarial losses and VAE regularization without any reconstruction losses improves the quality of the model for speech, it degrades the reconstruction quality for music. Semantic distillation is performed for the speech VAE similarly as in Mimi, with WavLM (Chen et al., 2021) as teacher. There is no semantic distillation for the music model and we let this for future work as semantic content is harder to define for music. The loss is:
$\mathcal{L}_{\text{VAE}}=\lambda_{\text{t}}\mathcal{L}_{\text{t}}(x,\hat{x})+\lambda_{\text{f}}\mathcal{L}_{\text{f}}(x,\hat{x})+\lambda_{\text{adv}}\mathcal{L}_{\text{adv}}(\hat{x})+\lambda_{\text{feat}}\mathcal{L}_{\text{feat}}(x,\hat{x})+\lambda_{\text{KL}}\mathcal{L}_{\text{KL}}$ (2)
where $\mathcal{L}_{\text{t}}$ and $\mathcal{L}_{\text{f}}$ are the temporal and frequential reconstruction losses, $\mathcal{L}_{\text{adv}}$ is the adversarial loss, $\mathcal{L}_{\text{feat}}$ is the feature matching loss and $\mathcal{L}_{\text{KL}}$ is the KL regularization applied to the VAE bottleneck.
4.2 Consistency-Based Continuous Audio Language Model (CALM)
Let $\bigl{(}\mathbf{x}^{1},\ldots,\mathbf{x}^{S}\bigr{)}$ denote the sequence of continuous latent vectors produced by a VAE encoder. As illustrated in Fig. 1, our model comprises three main components:
@@ 1. Causal Backbone Transformer with Noise Injection
We build on the MAR framework (Li et al., 2024) by employing a causal Transformer $T_{\text{long},\theta^{1}}$ to capture long-term dependencies. During training, each input latent $\mathbf{x}^{s}$ is replaced with a noised version $\tilde{\mathbf{x}}^{s}$ (see Sec. 4.3 and Ablation 5). We don’t perform any noise augmentation at inference time. Concretely, $\mathbf{z}_{\text{long}}^{s}\;=\;T_{\text{long},\theta^{1}}\bigl{(}\tilde{\mathbf{x}}^{1},\,\ldots,\,\tilde{\mathbf{x}}^{s-1}\bigr{)}$. Noise injection prevents error accumulation during inference, but as shown in Tab. 5 is insufficient alone for high-quality music generation.
2. Short-Context Transformer
To supply local, high-resolution context to the denoising head, we introduce a lightweight causal Transformer that attends the $K$ previous clean latents (we use $K=10$, $\sim$ 0.4s of music): $\mathbf{z}_{\text{short}}^{s}\;=\;T_{\text{short},\theta^{2}}\bigl{(}\mathbf{x}^{s-K},\,\ldots,\,\mathbf{x}^{s-1}\bigr{)}$. This short-context embedding $\mathbf{z}_{\text{short}}^{s}$ supplies fine-grained information potentially lost through noise injection in the backbone. We show in Tab. 7 that the value of K is not a decisive hyperparameter.
3. Consistency-Model Head
Finally, a small MLP-based consistency model $f_{\phi}$ is conditioned on the sum of the long-term and short-term features, $\mathbf{Z}^{s}=\mathbf{z}_{\text{long}}^{s}+\mathbf{z}_{\text{short}}^{s}$. At inference time, for 1-step, the next latent $\hat{\mathbf{x}}^{s}$ is sampled through: $\bm{\epsilon}\sim\mathcal{N}(0,I),t=1$, $\hat{\mathbf{x}}^{s}\;=\;f_{\phi}\bigl{(}\mathbf{x}^{s}_{1}=\bm{\epsilon},t=1,\mathbf{Z}^{s}\bigr{)}$.
In addition to consistency, we experiment with the TrigFlow (Lu & Song, 2025) formulation of flow-matching for the MLP. Although TrigFlow yields marginally higher fidelity, its inference cost makes it impractical for real-time use. This tradeoff is studied in Sec. 5.2, Tab. 4.
Together, these three components form a continuous autoregressive model that (i) leverages noise-robust long-term modeling, (ii) preserves local detail via short-context conditioning, and (iii) achieves rapid, high-fidelity latent sampling through consistency modeling.

### v3

@@ Abstract
Audio Language Models (ALM) have emerged as the dominant paradigm for speech and music generation by representing audio as sequences of discrete tokens. Yet, unlike text tokens, which are invertible, audio tokens are extracted from lossy codecs with a limited bitrate. As a consequence, increasing audio quality requires generating more tokens, which imposes a trade-off between fidelity and computational cost. We address this issue by studying Continuous Audio Language Models (CALM). These models instantiate a large Transformer backbone that produces a contextual embedding at every timestep. This sequential information then conditions an MLP that generates the next continuous frame of an audio VAE through consistency modeling. By avoiding lossy compression, CALM achieves higher quality at lower computational cost than their discrete counterpart. Experiments on speech and music demonstrate improved efficiency and fidelity over state-of-the-art discrete audio language models, facilitating lightweight, high-quality audio generation. Samples are available at iclr-continuous-audio-language-models.github.io. Finally, we release Pocket TTS, an open-source 100M-parameter text-to-speech model that can run faster than real time on a laptop CPU: github.com/kyutai-labs/pocket-tts.
@@ 1 Introduction
This motivates an alternative strategy: autoregressive modeling of continuous latents without quantization. Standard variational autoencoders (VAEs) are easier to train, are not affected by issues such as codebook collapse, and can reconstruct audio at higher fidelity for the same latent dimensionality. Pioneering work in the vision domain that autoregressively models continuous sequences includes GIVT (Tschannen et al., 2024) and MAR (Li et al., 2024), followed by some larger models in the image domain (Fan et al., 2025; Gu et al., 2025) and attempts in the audio domain (Turetzky et al., 2024; Pasini et al., 2024b; Jia et al., 2025; Yang et al., 2025). In MAR, the authors model the per-token probability distribution with a diffusion model (in the form of a small MLP) conditioned on a latent variable modeled by an autoregressive transformer backbone. SALAD (Turetzky et al., 2024) and DiTAR (Jia et al., 2025) adapt MAR-style diffusion heads for text-to-speech modeling, achieving better audio quality than discrete baselines. However, these works are limited to text-to-speech on small-scale and domain-specific datasets, leaving open the question of how well they can adapt to more complex tasks such as speech continuation (without text supervision) and richer audio domains such as music. We apply CALM to 4 tasks which are speech and music continuation as well as text-to-speech and text-to-music.
@@
We propose Continuous Audio Language Models (CALM) that predict sequences in the latent space of a VAE, bypassing the need for quantization. While we build on the MAR architecture where a transformer backbone uses $(\mathbf{x}^{1},\ldots,\mathbf{x}^{s-1})$ to predict an intermediate latent $\mathbf{z}^{s}$, which then conditions a head (diffusion model) that models $p(\mathbf{x}^{s}|\mathbf{z}^{s})$, we find that without further improvements, it fails to generate rich audio content and is slow to sample from. To overcome this, we introduce several key innovations:
@@
1. Improving quality and stability: To mitigate error accumulation during inference, we follow Pasini et al. (2024b) and, during training, inject noise into the long-term context $(\mathbf{x}^{1},\ldots,\mathbf{x}^{s-1})$. Additionally, we introduce a short-context transformer that summarizes recent clean latents, providing the sampling head with both coarse long-range context and fine-grained local information. 2. Diffusion-to-Consistency replacement: We replace the diffusion model with a continuous consistency model (Lu and Song, 2025) during training, significantly accelerating inference without compromising sample quality. This change reduces the inference time of the sampler head by a factor of up to $\times 20$ in our music experiments and $\times 12$ in our speech experiments compared to the use of a model that uses an RQ-Transformer as head. 3. Gaussian Temperature sampling: Temperature control is crucial for high-quality speech generation, yet consistency models lack a formal mechanism for temperature sampling. We present a heuristic that approximates temperature sampling in the consistency setup. 4. Head batch multiplier: Sampling multiple noise levels at training time for the same latent highly accelerates training for a small cost. 5. Latent Classifier Free Guidance: we apply Classifier Free Guidance to the latent variable conditioning the consistency head for the conditioned CALM. 6. Latent Distillation: once that we have chosen a latent CFG coefficient, we can distill the CFG computation of the backbone into a student backbone while keeping the same sampling head (MLP), hence dividing the batch size by 2 at inference time. As well, we can apply the distillation to a much smaller student backbone.
Finally, by using these innovations, we introduce Pocket TTS which is a 100M-parameters text-to-speech model that can run faster than real-time on a laptop CPU. We detail the results of Pocket TTS in Sec.F and in the following technical report: kyutai.org/pocket-tts-technical-report.
@@ 2 Related Work
Autoregressive audio language models. Early autoregressive audio models operated on raw waveforms such as WaveNet (van den Oord et al., 2016), learned discrete codes via VQ-VAE, such as in Jukebox (Dhariwal et al., 2020) or continuous word-sized audio tokens (Algayres et al., 2023). The advent of neural audio codecs (Zeghidour et al., 2021; Défossez et al., 2024a) enabled high-quality vector-quantized tokenizers for general audio. These, in turn, powered discrete-token audio LMs: AudioLM (Borsos et al., 2023) for unconditioned audio generation while AudioGen (Kreuk et al., 2023), MusicLM (Agostinelli et al., 2023), and MusicGen (Copet et al., 2023) apply similar methods for text-to-audio and text-to-music generation. In speech, autoregressive modeling of RVQ codes has been used for text-to-speech generation (Wang et al., 2023; Kharitonov et al., 2023) as well as for spoken dialogue (Défossez et al., 2024b) or translation (Labiausse et al., 2025). However, all of these systems rely on lossy quantization, which inevitably degrades audio fidelity unless a significant compute budget is spent to generate a deep hierarchy of RVQ tokens. This is unlike CALM, which predicts continuous embeddings in one pass, providing a better quality-computation trade-off.
@@
Continuous autoregressive models. In GIVT, Tschannen et al. (2024) use a transformer to autoregressively model the latent space of a VAE trained on images, parameterize a Gaussian Mixture Model, and train it with a cross-entropy loss. The authors of MAR (Li et al., 2024) obtain better results by replacing the Gaussian mixture model with a diffusion-model enabling the approximation of more diverse distributions. In MAR, a large transformer backbone predicts a continuous embedding $\mathbf{z}^{s}$ given $(\mathbf{x}^{1},\ldots,\mathbf{x}^{s-1})$, which then conditions a small MLP diffusion network that models the probability distribution $p(\mathbf{x}^{s}|\mathbf{z}^{s})$ of the next latent. This eliminates the need for discrete tokenizers, but at the cost of slow sampling: MAR typically needs hundreds of denoising steps per token. Hang et al. (2025) aims to speed up MAR by replacing the diffusion head with a shortcut head (Frans et al., 2025) for few-step sampling. Shortcut models combine a diffusion loss and a self-consistency loss in order to accelerate the diffusion process. They reduce the number of diffusion steps from 100 to 8 with a similar image quality. Remarkably, CALM achieves a quality comparable to the best discrete models with only one step of consistency modeling.
@@
In audio, the approach of MAR has been adapted for the task of Text-to-Speech (TTS) synthesis. SALAD (Turetzky et al., 2024) introduces a zero-shot TTS model that operates on continuous speech representations using a per-token latent diffusion process. By leveraging semantic tokens for contextual information and determining synthesis stopping points, SALAD achieves improved intelligibility and audio quality without relying on quantization. Similarly, DiTAR (Jia et al., 2025) presents a patch-based autoregressive framework combining a language model with a diffusion transformer for speech generation. This approach models dependencies between aggregated local patches of continuous tokens, using a causal language model to produce embeddings, which, along with previous patches, serve as inputs to a bidirectional diffusion transformer that predicts the next patch. The authors observe that providing local context through patching was determinant to improve their model, which corroborates what we observe with the introduction of a short context transformer into our model. In Pasini et al. (2024b), the authors apply the MAR framework to music generation using a relatively small dataset comprising 20,000 single-instrument stems, training their model on 10-second excerpts on top of their continuous compression model Music2Latent (Pasini et al., 2024a). They introduce a method for noise augmentation of the data that allows the model to avoid error accumulation. However, we notice that scaling to a more complex and diverse dataset consisting of full musical pieces makes their approach struggle to maintain high-quality generation on longer sequences. To address these challenges, we propose novel strategies that enhance generation quality while also improving inference efficiency. Finally, Music2Latent2 (Pasini et al., 2025) explores combining autoregressive and consistency modeling but for compression. IMPACT (Huang et al., 2025) explores MAR decoding for text-to-audio generation and sets a new standard for short latency models on the AudioCaps (Kim et al., 2019) benchmark. More recently, MingUni-Audio (Yan et al., 2025) shows that continuous speech language models can scale to 20B Mixture of Experts models with 3B active parameters.
Some other autoregressive speech models work in the continuous domain thanks to spectral representations such as MELLE (Meng et al., 2024) for TTS and Flow-Omni (Yuan et al., 2024) for speech-to-speech conversation.
@@ 3 Background
This approach enables efficient parallel modeling of RVQ sequences, allowing all codebooks corresponding to a single timestep to be generated simultaneously without introducing any delay. However, a key limitation lies in the use of the RQ-Transformer, which is computationally intensive; as its resource requirements scale with the number of RVQ, or even with the square of the number of RVQ, if the attention dominates the computation cost.
@@ 3.2 Consistency Models
By using $T=\frac{\pi}{2}$ and $\alpha_{t}=\cos(t),\sigma_{t}=\sin(t)$, Lu and Song (2025) derive the following continuous-time consistency loss where $w_{\psi}(t)$ is an adaptive weighting function:
@@
Lagrangian Self-Distillation. We also explore a new 1-step flow-matching method named Latent Self-Distillation (LSD) that has been introduced in Boffi et al. (2025). In this paper, the authors unify most 1-step methods into 2 categories and develops a third one (LSD) which appears to be more stable at training. See Sec. A for the equations.
@@ 4.1 Our VAE-GAN
While training the model with adversarial losses and VAE regularization without any reconstruction losses improves the quality of the model for speech, it degrades the reconstruction quality for music. Semantic distillation is performed for the speech VAE similarly as in Mimi, with WavLM (Chen et al., 2021b) as teacher. There is no semantic distillation for the music model and we let this for future work as semantic content is harder to define for music. The loss is:
$\mathcal{L}_{\text{VAE}}=\lambda_{\text{t}}\mathcal{L}_{\text{t}}(x,\hat{x})+\lambda_{\text{f}}\mathcal{L}_{\text{f}}(x,\hat{x})+\lambda_{\text{adv}}\mathcal{L}_{\text{adv}}(\hat{x})+\lambda_{\text{feat}}\mathcal{L}_{\text{feat}}(x,\hat{x})+\lambda_{\text{KL}}\mathcal{L}_{\text{KL}}+\lambda_{\text{distill}}\mathcal{L}_{\text{distill}}$ (2)
where $\mathcal{L}_{\text{t}}$ and $\mathcal{L}_{\text{f}}$ are the temporal and frequential reconstruction losses, $\mathcal{L}_{\text{adv}}$ is the adversarial loss, $\mathcal{L}_{\text{feat}}$ is the feature matching loss, $\mathcal{L}_{\text{KL}}$ is the KL regularization applied to the VAE bottleneck, and $\mathcal{L}_{\text{distill}}$ is the WavLM distillation loss applied for the speech VAE.
4.2 Continuous Audio Language Model (CALM) architecture
Let $\bigl(\mathbf{x}^{1},\ldots,\mathbf{x}^{S}\bigr)$ denote the sequence of continuous latent vectors produced by a VAE encoder. As illustrated in Fig. 1, our model comprises three main components. The motivations behind these design choices are described in Sec. 4.3 and the ablation study in Tab. 5 justifies these choices.
@@ 1. Causal Backbone Transformer with Noise Injection
We build on the MAR framework (Li et al., 2024) by employing a causal Transformer $T_{\text{long},\theta^{1}}$ to capture long‑term dependencies. However, during preliminary experiments, we realized that music generation models with the MAR framework were generating poor quality audio and diverging quickly at inference because they were not robust to error accumulation. Pasini et al. (2024b) introduced a noise augmentation trick at training time in order to tackle this problem. Given a sequence $(\mathbf{x}^{1},\mathbf{x}^{2},\ldots,\mathbf{x}^{S})$, they sample $k_{s}\sim\mathcal{U}(0,1)$ and $\bm{\epsilon}_{s}\sim\mathcal{N}(0,\mathbf{I})$ for every $s\in\{1,\ldots,S\}$ and use a noised input to the backbone $\tilde{\mathbf{x}}^{s}=k_{s}\bm{\epsilon}_{s}+(1-k_{s})\mathbf{x}^{s}$ for every $s$. Early experiments showed that preserving the variance of $\tilde{\mathbf{x}}^{s}$ improved quality, so that we use instead $\tilde{\mathbf{x}}^{s}=\sqrt{k_{s}}\bm{\epsilon}_{s}+\sqrt{1-k_{s}}\mathbf{x}^{s}$. We don’t perform any noise augmentation at inference time. Thus, $\mathbf{z}_{\text{long}}^{s}\;=\;T_{\text{long},\theta^{1}}\bigl(\tilde{\mathbf{x}}^{1},\,\ldots,\,\tilde{\mathbf{x}}^{s-1}\bigr)$. Noise injection prevents error accumulation during inference, but as shown in Tab. 5 is insufficient alone for high‑quality music generation.
2. Short‑Context Transformer
To supply local, high‑resolution context to the denoising head, we introduce a lightweight causal Transformer that attends the $K$ previous clean latents (we use $K=10$, $\sim$ 0.4s of music): $\mathbf{z}_{\text{short}}^{s}\;=\;T_{\text{short},\theta^{2}}\bigl(\mathbf{x}^{s-K},\,\ldots,\,\mathbf{x}^{s-1}\bigr)$. This short‑context embedding $\mathbf{z}_{\text{short}}^{s}$ supplies fine‑grained information potentially lost through noise injection in the backbone. We show in Sec. E.2 that the value of K is not a decisive hyperparameter but Tab. 5 indicates that the short-context transformer is crucial to good quality generation.
3. Consistency‑Model Head
Finally, a small MLP-based consistency model $f_{\phi}$ is conditioned on the sum of the long‑term and short‑term features, $\mathbf{Z}^{s}=\mathbf{z}_{\text{long}}^{s}+\mathbf{z}_{\text{short}}^{s}$. At inference time, for 1-step generation, the next latent $\hat{\mathbf{x}}^{s}$ is sampled through: $\bm{\epsilon}\sim\mathcal{N}(0,I),t=1$, $\hat{\mathbf{x}}^{s}\;=\;f_{\phi}\bigl(\mathbf{x}^{s}_{1}=\bm{\epsilon},t=1,\mathbf{Z}^{s}\bigr)$.
In addition to consistency, we experiment with the TrigFlow (Lu and Song, 2025) formulation of flow‑matching for the MLP. Although TrigFlow yields marginally higher fidelity, its inference cost makes it impractical for real‑time use. While Tab. 4 shows it, this tradeoff is studied in Sec. E.3.
Together, these three components form a continuous autoregressive model that (i) leverages noise‑robust long‑term modeling, (ii) preserves local detail via short‑context conditioning, and (iii) achieves rapid, high‑fidelity latent sampling through consistency modeling.